OpenStudy (anonymous):

Solve the equation for y. 6x-3y=1

OpenStudy (australopithecus):

-6x + 6x -3y = 1 - 6x

OpenStudy (australopithecus):

-3y = 1 - 6x

OpenStudy (australopithecus):

\[\frac{-3y}{-3} = \frac{1-6x}{-3}\] \[y = \frac{1-6x}{-3}\]

OpenStudy (australopithecus):

remember -3/-3 = 1 and 1(y) = y

OpenStudy (lopus):